Calculus I, Fall 2015 (http://www.math.nthu.edu.tw/˜ wangwc/)

Homework Assignment for Week 16

1. Section 8.1: Problems 5, 11, 15, 23, 25, 27, 29, 33, 35, 65, 71, 73.

2. Section 8.2: Problems 11, 17, 27, 28, 33, 34, 35, 37, 45, 57, 63, 65, 67.

3. Section 8.2: Derive formula 75 76 of ’A brief table of integrals’ near the end of the textbook.

4. Section 8.3: Problems 3, 5, 13, 35, 37, 45, 47, 48, 54, 57.

5. Section 8.4: Problems 1, 3, 21, 23, 27, 29, 32, 35, 37, 39, 41, 43, 45, 47, 49.

6. Section 8.4:

7. Chap 8: Problems 41, 43, 45, 47, 49 (p 512).

8. Chap 8: (Important) As time permits, do as many as you can in odd number problems 69-115, p 509. They contain all the integration techniques and you have to figure out which one(s) to use for each problem.